OpenStudy (anonymous):

Use the Binomial Theorem to find the binomial expansion of the expression. (x-3)^5

OpenStudy (anonymous):

please Help me Somebody.... will give medal ..

OpenStudy (anonymous):

I think you can find it ! Help this : (x-3)(x-3)(x-3)(x-3)(x-3)= ? Use from Broadcast :)

OpenStudy (anonymous):

wait what do you Mean use from broadcast? ;o

OpenStudy (anonymous):

(x-3)(x-3)=x * x + x* -3 -3 * x -3*-3 k ?:)